### v2

@@ Appendix B Proof of Section 3
$=\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
Clearly, $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)=-\infty$ if $\mathbb{P}\left(\cap_{i=1}^{n}B_{i}\right)=0$. Next, we consider the case $\mathbb{P}\left(\cap_{i=1}^{n}B_{i}\right)>0$. Note that $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ and some $(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$ is equivalent to $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},\cap_{i=1}^{n}B_{i}\right)\leqslant\sum_{i=1}^{n}\mathbb{P}(B_{i})\Lambda_{i}(y_{i})$ for some $(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$. Using this conclusion and the attainability of $\Lambda^{\diamond}$, we have
$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)$ $=\inf\left\{x:x=\sum_{i=1}^{n}y_{i},\mathbb{P}\left(X>x,\cap_{i=1}^{n}B_{i}\right)\leqslant\sum_{i=1}^{n}\mathbb{P}(B_{i})\Lambda_{i}(y_{i})\right\}$
$=\inf\left\{x:x=\sum_{i=1}^{n}y_{i},\mathbb{Q}\left(X>x\right)\leqslant\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\right\}$
@@
$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(X>x,A^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{Q}_{2}\left(X>x,A\right)\leqslant\Lambda_{2}(x-y),~{}\text{for some}~{}y\in\mathbb{R},A\in\mathcal{F}\right\}$
$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(X>x,A^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,A\}})\leqslant\Lambda_{2}(x-y),~{}\text{for some}~{}y\in\mathbb{R},A\in\mathcal{F}\right\}.$
Let $f(t)=\mathbb{Q}_{1}(X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant t),~{}t\in[0,1]$. Clearly, $f$ is an increasing and continuous function. Let $f^{-1}$ denote the left-quantile of $f$. Define $A_{x,t}=\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant f^{-1}(t)\}$ for $t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$. Let $t_{0}=\mathbb{Q}_{1}(X>x,A)$. Then it follows that $\mathbb{Q}_{1}(X>x,A_{x,t_{0}})=\mathbb{Q}_{1}(X>x,A)$ and
$\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,A_{x,t_{0}}\}})=\mathbb{E}^{\mathbb{Q}_{1}}(F_{\eta}^{\mathbb{Q}_{1},-1}(U_{\eta}^{\mathbb{Q}_{1}})\mathds{1}_{\{X>x,A_{x,t_{0}}\}})\leqslant\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,A\}}).$
@@
$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(X>x,A_{x,t}^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,A_{x,t}\}})\leqslant\Lambda_{2}(x-y),~{}\text{for some}~{}y\in\mathbb{R},t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]\right\}$
@@
$\mathbb{Q}_{1}(X_{1}>y^{*})$ $=\mathbb{Q}_{1}(X>x^{*},\Omega\setminus A^{*})=\mathbb{Q}_{1}(X>x^{*})-\mathbb{Q}_{1}(X>x^{*},A^{*})$
@@
$\mathbb{Q}_{2}(X_{2}>x^{*}-y^{*})$ $=\mathbb{Q}_{2}(X>x^{*},A^{*})=\mathbb{E}^{\mathbb{Q}_{1}}\left(\eta\mathds{1}_{\{X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant z^{*}\}}\right)$
@@
Appendix C Proof of Section 4
In this section, we provide all the proofs of results in Section 4.
@@ Appendix C Proof of Section 4
We next show the inverse inequality. For any $X_{1}\in\mathcal{X}$, let $x_{1}=\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})$. Note that $X_{1}=F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})$ a.s. under $\mathbb{Q}_{1}$. Define $Y_{1}=x_{1}\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}+(X+m)\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}}$ with $m>(x_{1}-\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{1}}X)\vee(\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}|X|+\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{2}}|X_{1}|+|x_{1}|)$, where $\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{i}}$ and $\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{i}}$ represent the $\mathrm{ess\mbox{-}inf}$ and $\mathrm{ess\mbox{-}sup}$ under $\mathbb{Q}_{i}$, $i=1,2$. Note that $U_{X_{1}}^{\mathbb{Q}_{1}}$ can be chosen such that $\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant t\}\subset\{X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(t)\}$ for all $t\in(0,1)$. By the definition of $\Lambda\mathrm{VaR}$ and the right continuity of $\Lambda$, we have $F_{X_{1}}^{\mathbb{Q}_{1}}(x_{1})\geqslant 1-\Lambda(x_{1})$, which implies $x_{1}\geqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(1-\Lambda(x_{1}))$. Hence we have $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(Y_{1})=x_{1}$, and $X_{1}\leqslant Y_{1}$ for all $w\in\Omega$. By monotonicity of $\rho$, we have
@@
We next show that the existence of the optimal allocation implies the existence of the minimizer of (5). Suppose there exists $X_{1}\in\mathcal{X}$ such that $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})=\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)$. Following the same argument as above to show the inverse inequality, let $x_{1}=\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})$ and $Y_{1}=x_{1}\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}+(X+m)\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}}$ with $m>(x_{1}-\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{1}}X)\vee(\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}|X|+\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{2}}|X_{1}|+|x_{1}|)$ such that $X_{1}\leqslant Y_{1}$ for all $w\in\Omega$. We have
@@
For $y<\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{2}}(X-x)$, we have $(X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\geqslant y$ a.s. under $\mathbb{Q}_{2}$. It follows from the monotonicity and law-invariance of $\rho$ that $\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\geqslant\rho^{\mathbb{Q}_{2}}(y)$. Using the monotonicity of $\rho$ again, we have $\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}\geqslant\inf_{y\in\mathbb{R}}\left\{x+\rho^{\mathbb{Q}_{2}}(y)\right\}.$ Consequently,
@@
Next, we consider the case $\mathbb{P}(B_{1}\cap B_{2}^{c})<(1-\Lambda(x))\mathbb{P}(B_{1})$. Let $y<\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{2}}(X-x)$ and $A_{x}=(B_{1}\cap B_{2}^{c})\cup(B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})$ with $\alpha_{x}$ satisfying $\mathbb{P}(B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})=(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})$. For any $B\in\mathcal{F}$ with $\mathbb{Q}_{1}(B)=1-\Lambda(x)$ and $z\geqslant y$, we have
$\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)=\frac{\mathbb{P}(X\leqslant z+x,B\cap B_{2})+\mathbb{P}(B^{c}\cap B_{2})}{\mathbb{P}(B_{2})}=1-\frac{\mathbb{P}(X>z+x,B\cap B_{2})}{\mathbb{P}(B_{2})}.$
@@
$\mathbb{P}(X>z+x,A_{x}\cap B_{2})=\mathbb{P}(X>z+x,B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})\leqslant\mathbb{P}(X>z+x,B_{1}\cap B_{2}\cap B_{3})$
for any $B_{3}\in\mathcal{F}$ such that $\mathbb{P}(B_{1}\cap B_{2}\cap B_{3})\geqslant(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})$. Note that $\mathbb{Q}_{1}(B)=1-\Lambda(x)$ implies $\mathbb{P}(B_{1}\cap B_{2}\cap B)\geqslant(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})$. Hence we have $\mathbb{P}(X>z+x,A_{x}\cap B_{2})\leqslant\mathbb{P}(X>z+x,B_{1}\cap B_{2}\cap B)\leqslant\mathbb{P}(X>z+x,B\cap B_{2})$. This implies
$\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)\leqslant\mathbb{Q}_{2}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}}\leqslant z)$
for $z\geqslant y$. For $z<y$, clearly,

### v3

@@ Appendix A Proof of all results in Sections 2-5
$=\inf\left\{\sum_{j=1}^{n}y_{j}:\mathbb{P}\left(\left\{X>\sum_{j=1}^{n}y_{j}\right\}\cap\left(\cap_{j=1}^{n}B_{j}\right)\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~\text{for some}~(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
Clearly, $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)=$-$\infty$ if $\mathbb{P}\left(\cap_{i=1}^{n}B_{i}\right)=0$. Next, we consider the case $\mathbb{P}\left(\cap_{i=1}^{n}B_{i}\right)>0$. Note that $\mathbb{P}\left(\{X>\sum_{j=1}^{n}y_{j}\}\cap\left(\cap_{j=1}^{n}B_{j}\right)\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ and some $(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$ is equivalent to $\mathbb{P}\left(\{X>\sum_{j=1}^{n}y_{j}\}\cap(\cap_{j=1}^{n}B_{j})\right)\leqslant\sum_{i=1}^{n}\mathbb{P}(B_{i})\Lambda_{i}(y_{i})$ for some $(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$. Using this conclusion and the attainability of $\Lambda^{\diamond}$, we have
$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)$ $=\inf\left\{x:x=\sum_{i=1}^{n}y_{i},\mathbb{P}\left(\{X>x\}\cap(\cap_{j=1}^{n}B_{j})\right)\leqslant\sum_{i=1}^{n}\mathbb{P}(B_{i})\Lambda_{i}(y_{i})\right\}$
$=\inf\left\{x:x=\sum_{i=1}^{n}y_{i},\mathbb{Q}\left(X>x\right)\leqslant\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})}\right\}$
@@
$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(\{X>x\}\cap A^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{Q}_{2}\left(\{X>x\}\cap A\right)\leqslant\Lambda_{2}(x-y),~\text{for some}~y\in\mathbb{R},A\in\mathcal{F}\right\}$
$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(\{X>x\}\cap A^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x\}\cap A})+\mathbb{Q}_{2}(\{X>x\}\cap A\cap N)\leqslant\Lambda_{2}(x-y),\right$.
$~~~~~~~~~~~~\left.\text{for some}~y\in\mathbb{R},A\in\mathcal{F}\right\}.$
Let $A^{\prime}=A\cap N^{c}$. Using the fact $\mathbb{Q}_{1}(N)=0$, we have $\mathbb{Q}_{1}\left(\{X>x\}\cap(A^{\prime})^{c}\right)=\mathbb{Q}_{1}\left(\{X>x\}\cap A^{c}\right)$, $\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x\}\cap A^{\prime}})=\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x\}\cap A})$ and $\mathbb{Q}_{2}(\{X>x\}\cap A^{\prime}\cap N)=0$. Hence, the inf-convolution can be simplified as
$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(\{X>x\}\cap A^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x\}\cap A})\leqslant\Lambda_{2}(x-y),~\text{for some}~y\in\mathbb{R},A\in\mathcal{F}\right\}.$
Let $f(t)=\mathbb{Q}_{1}(\{X>x\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant t\}),~t\in[0,1]$. Clearly, $f$ is an increasing and continuous function. Let $f^{-1}$ denote the left-quantile of $f$. Define $A_{x,t}=\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant f^{-1}(t)\}$ for $t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$. Let $t_{0}=\mathbb{Q}_{1}(\{X>x\}\cap A)$. Then it follows that $\mathbb{Q}_{1}(\{X>x\}\cap A_{x,t_{0}})=\mathbb{Q}_{1}(\{X>x\}\cap A)$ and
$\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x\}\cap A_{x,t_{0}}})=\mathbb{E}^{\mathbb{Q}_{1}}(F_{\eta}^{\mathbb{Q}_{1},-1}(U_{\eta}^{\mathbb{Q}_{1}})\mathds{1}_{\{X>x\}\cap A_{x,t_{0}}})\leqslant\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x\}\cap A}).$
@@
$=\inf\left\{x\in\mathbb{R}:\mathbb{Q}_{1}\left(\{X>x\}\cap A_{x,t}^{c}\right)\leqslant\Lambda_{1}(y),\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x\}\cap A_{x,t}})\leqslant\Lambda_{2}(x-y),\right$.
$~~~~~~~~~~~~\left.\text{for some}~y\in\mathbb{R},t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]\right\}$
@@
$\mathbb{Q}_{1}(X_{1}>y^{*})$ $=\mathbb{Q}_{1}(\{X>x^{*}\}\cap(\Omega\setminus A^{*}))=\mathbb{Q}_{1}(X>x^{*})-\mathbb{Q}_{1}(\{X>x^{*}\}\cap A^{*})$
@@
$\mathbb{Q}_{2}(X_{2}>x^{*}-y^{*})$ $=\mathbb{Q}_{2}(\{X>x^{*}\}\cap A^{*})=\mathbb{E}^{\mathbb{Q}_{1}}\left(\eta\mathds{1}_{\{X>x^{*}\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant z^{*}\}}\right)$
@@
Proof of Corollary 2. Note that $\mathbb{Q}_{2}(N)\geqslant 1-\lambda_{2}^{-}$ implies $g_{x}(t)\leqslant 1-\mathbb{Q}_{2}(N)\leqslant\lambda_{2}^{-}\leqslant\Lambda_{2}(x-y)$ for all $x,y\in\mathbb{R}$. Hence, the inf-convolution is -$\infty$. ∎
@@
We next show the inverse inequality. For any $X_{1}\in\mathcal{X}$, let $x_{1}=\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})$. Note that $X_{1}=F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})$ a.s. under $\mathbb{Q}_{1}$. Define $Y_{1}=x_{1}\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}+(X+m)\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}}$ with $m>m_{0}$. Note that $U_{X_{1}}^{\mathbb{Q}_{1}}$ can be chosen such that $\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant t\}\subset\{X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(t)\}$ for all $t\in(0,1)$. By the definition of $\Lambda\mathrm{VaR}$ and the right continuity of $\Lambda$, we have $F_{X_{1}}^{\mathbb{Q}_{1}}(x_{1})\geqslant 1-\Lambda(x_{1})$, which implies $x_{1}\geqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(1-\Lambda(x_{1}))$. Hence, $X_{1}\leqslant x_{1}$ for all $w\in\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}$.
For the case $\mathcal{X}=L^{\infty}$, let $m_{0}=(x_{1}-\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{1}}X)\vee(\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}|X|+\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{2}}|X_{1}|+|x_{1}|)$, where $\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{i}}$ and $\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{i}}$ represent the $\mathrm{ess\mbox{-}inf}$ and $\mathrm{ess\mbox{-}sup}$ under $\mathbb{Q}_{i}$, $i=1,2$. Then we have $X_{1}\leqslant Y_{1}$ a.s. under $\mathbb{Q}_{2}$, which further implies $X-X_{1}\geqslant X-Y_{1}$ a.s. under $\mathbb{Q}_{2}$. By the monotonicity of $\rho$, we have
@@
$=x_{1}+\rho^{\mathbb{Q}_{2}}\left((X-x_{1})\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}-m\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}}\right)$
For the case that $\rho^{\mathbb{Q}_{2}}$ is an $\varepsilon$-tail risk measure, let $m_{0}=-\mathrm{VaR}_{\varepsilon}^{\mathbb{Q}_{2}}(X-X_{1})$. Then we have for $x\geqslant\mathrm{VaR}_{\varepsilon}^{\mathbb{Q}_{2}}(X-X_{1})$
$\mathbb{Q}_{2}(X-X_{1}\leqslant x)$
$=\mathbb{Q}_{2}(\{X-X_{1}\leqslant x\}\cap\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\})+\mathbb{Q}_{2}(\{X-X_{1}\leqslant x\}\cap\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\})$
$\leqslant\mathbb{Q}_{2}(\{X-x_{1}\leqslant x\}\cap\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\})+\mathbb{Q}_{2}(U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1}))$
$=\mathbb{Q}_{2}(X-Y_{1}\leqslant x),$
which implies
$\frac{(F_{X-X_{1}}^{\mathbb{Q}_{2}}(x)-(1-\varepsilon))_{+}}{\varepsilon}\leqslant\frac{(F_{X-Y_{1}}^{\mathbb{Q}_{2}}(x)-(1-\varepsilon))_{+}}{\varepsilon},~x\in\mathbb{R}.$
By the monotonicity of $\rho$ and the fact $\rho$ is an $\varepsilon$-tail risk measure, we have
$=x_{1}+\rho^{\mathbb{Q}_{2}}\left((X-x_{1})\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}-m\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}}\right)$
@@
We next show that the existence of the optimal allocation implies the existence of the minimizer of (6). Suppose there exists $X_{1}\in\mathcal{X}$ such that $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})=\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho^{\mathbb{Q}_{2}}(X)$. Following the same argument as above to show the inverse inequality, let $x_{1}=\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}(X_{1})$ and $Y_{1}=x_{1}\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant 1-\Lambda(x_{1})\}}+(X+m)\mathds{1}_{\{U_{X_{1}}^{\mathbb{Q}_{1}}>1-\Lambda(x_{1})\}}$ with some $m>m_{0}$ such that $X_{1}\leqslant Y_{1}$ a.s. under $\mathbb{Q}_{2}$ if $\mathcal{X}=L^{\infty}$, and $(F_{X-X_{1}}^{\mathbb{Q}_{2}}(x)-(1-\varepsilon))_{+}\leqslant(F_{X-Y_{1}}^{\mathbb{Q}_{2}}(x)-(1-\varepsilon))_{+},~x\in\mathbb{R}$ if $\mathcal{X}$ is unbounded and $\rho^{\mathbb{Q}_{2}}$ is an $\varepsilon$-tail risk measure. We have
@@
If $\mathcal{X}=L^{\infty}$, for $y<\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{2}}(X-x)$, we have $(X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\geqslant y$ a.s. under $\mathbb{Q}_{2}$. It follows from the monotonicity and law-invariance of $\rho$ that $\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\geqslant\rho^{\mathbb{Q}_{2}}(y)$. Using the monotonicity of $\rho$ again, we have
$\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}\geqslant\inf_{y\in\mathbb{R}}\left\{x+\rho^{\mathbb{Q}_{2}}(y)\right\}.$
If $\rho^{\mathbb{Q}_{2}}$ is an $\varepsilon$-tail risk measure, for $y<\mathrm{VaR}_{\varepsilon}^{\mathbb{Q}_{2}}(X-x)$, we have for $z\geqslant y$
$\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)=\mathbb{Q}_{2}(\{X-x\leqslant z\}\cap B)+\mathbb{Q}_{2}(B^{c})\leqslant 1=\mathbb{Q}_{2}(y\leqslant z).$
Moreover, $\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}<y)=\mathbb{Q}_{2}(\{X-x<y\}\cap B)<1-\varepsilon$. Hence, $\mathrm{VaR}_{\varepsilon}^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\geqslant y$ and
$\frac{(\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)-(1-\varepsilon))_{+}}{\varepsilon}\leqslant\frac{(\mathds{1}_{[y,\infty)]}(z)-(1-\varepsilon))_{+}}{\varepsilon},~z\in\mathbb{R}.$
Using the monotonicity of $\rho$ and the fact that $\rho$ is an $\varepsilon$-tail risk measure under $\mathbb{Q}_{2}$, we have
$\inf_{y\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\right\}\geqslant\inf_{y\in\mathbb{R}}\left\{x+\rho^{\mathbb{Q}_{2}}(y)\right\}.$
Consequently,
@@
Next, we consider the case $\mathbb{P}(B_{1}\cap B_{2}^{c})<(1-\Lambda(x))\mathbb{P}(B_{1})$. Let $A_{x}=(B_{1}\cap B_{2}^{c})\cup(B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})$ with $\alpha_{x}$ satisfying $\mathbb{P}(B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})=(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})$. For any $B\in\mathcal{F}$ with $\mathbb{Q}_{1}(B)=1-\Lambda(x)$ and $z\geqslant y$, we have
$\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)$ $=\frac{\mathbb{P}(\{X\leqslant z+x\}\cap B\cap B_{2})+\mathbb{P}(B^{c}\cap B_{2})}{\mathbb{P}(B_{2})}$
$=1-\frac{\mathbb{P}(\{X>z+x\}\cap B\cap B_{2})}{\mathbb{P}(B_{2})}.$
@@
$\mathbb{P}(\{X>z+x\}\cap A_{x}\cap B_{2})=\mathbb{P}(\{X>z+x\}\cap B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})\leqslant\mathbb{P}(\{X>z+x\}\cap B_{1}\cap B_{2}\cap B_{3})$
for any $B_{3}\in\mathcal{F}$ such that $\mathbb{P}(B_{1}\cap B_{2}\cap B_{3})\geqslant(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})$. Note that $\mathbb{Q}_{1}(B)=1-\Lambda(x)$ implies $\mathbb{P}(B_{1}\cap B_{2}\cap B)\geqslant(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})$. Hence we have $\mathbb{P}(\{X>z+x\}\cap A_{x}\cap B_{2})\leqslant\mathbb{P}(\{X>z+x\}\cap B_{1}\cap B_{2}\cap B)\leqslant\mathbb{P}(\{X>z+x\}\cap B\cap B_{2})$. This implies
$\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)\leqslant\mathbb{Q}_{2}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}}\leqslant z),~z\geqslant y.$
For $\mathcal{X}=L^{\infty}$, let $y<\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{2}}(X-x)$. Then for $z<y$, clearly,